OpenStudy (anonymous):

Three consecutive odd integers have a sum of 27. Find the integers.

OpenStudy (studygurl14):

N + (N + 2) + (N + 4) = 27 3N + 6 = 27 3N = 21 N = 7' 7, 9, 11

OpenStudy (the_fizicx99):

3x + 3 = 27 -3 -3 3x = 24 /3 /3 x = 8 8 + 9 + 10 = 27

OpenStudy (anonymous):

Thank you both!!!

OpenStudy (studygurl14):

@tHe_FiZiCx99 , the integers are all ODD!

OpenStudy (the_fizicx99):

Study girl its x + (x+1) +( x+2)

OpenStudy (the_fizicx99):

"Odd", Yea, mine are even. Well only 8 and 10 I believe is even

OpenStudy (studygurl14):

No it isn't. that is the formula if the question said :three consecutive integers, not three consecutive ODD integers. All the integers have to be odd. since every other number is odd, then it is x, x + 2, x + 4, etc.

OpenStudy (anonymous):

Can you guys help me with one more?

OpenStudy (studygurl14):

sure

OpenStudy (anonymous):

Btw, studygirl was correct

OpenStudy (anonymous):

Two consecutive integers have a sum of 159. Find the integers.

OpenStudy (anonymous):

@StudyGurl14 ?

OpenStudy (anonymous):

nvmind I'll post it as a new question

OpenStudy (studygurl14):

@getoffmyback N + (N + 2) = 159 2N + 2 = 159 2N = 157 N = not possible. I think you wrote the problem down wrong...

OpenStudy (amriju):

@StudyGurl14 ..its n+(n+1)..dont u think..consecutive integers... N+(N+1)=159..2N+1=159...2N=158...N=59..N+1=60

OpenStudy (studygurl14):

oh, sorry, no wonder...lolthanks @amriju